# Resource-Constrained Federated Continual Learning: What Does Matter?

### v1

@@ III-A Experiment Setup
TABLE III: Experimental Details. Analysis of various considered settings of different datasets in the experiments section.
@@
Image Size 3$\times$32$\times$32 3$\times$32$\times$32 3$\times$64$\times$64 1$\times$28$\times$28 3$\times$300$\times$300 3$\times$300$\times$300
@@
III-B Overview Performance under Resource Constraints
In this section, we systematically investigate the influence of resource constraints on existing FCL methods and typical techniques studied in the FCL literature, namely Sample Caching, Data Synthesis, Knowledge Distillation, and Network Extension.
With Sufficient Resources. In Table II(a), our initial assessment of the FCL methods under sufficient resource conditions reveals a promising performance across different datasets. Notably, both FedCIL and AF-FCL methods excel, surpassing most established baselines on CIFAR-10 and Digit-10. This superior performance underscores the effectiveness of their generative models, which can effectively train on relatively simple datasets and generate high-quality samples for data replay. However, as we delve into more complex and larger-scale datasets like CIFAR-100 and Tiny-ImageNet, we observe a significant drop in performance for both FedCIL and AF-FCL. This decline highlights the limitations of their generative models when faced with greater data diversity and complexity. In the context of federated domain-incremental learning, where the data distribution changes over time, FCIL demonstrates an interesting behavior. When there are no new incremental sample classes, it reverts to the FedAvg algorithm, a widely used federated learning baseline. On the other hand, Re-Fed consistently achieves reasonable performance across all datasets and scenarios. Its synergistic replay strategy, which combines both generated and real data, seems to provide a robust foundation for FCL.
With Limited Resources. The experiments conducted under extremely limited resource conditions starkly contrast to the results obtained with sufficient resources. By simultaneously restricting three key types of training resources, we observe a significant decline in the performance of all methods. This underscores the critical importance of these resources for the effectiveness of federated continual learning techniques. To better understand the impact of these resource constraints, we conducted experiments from two perspectives. First, we analyzed how each resource limitation affects the performance of the various methods. This allowed us to identify the most sensitive components and bottlenecks that hinder the effectiveness of the FCL techniques under limited conditions. Second, we explored the interactions between different resource constraints and how they compound the challenges faced by the FCL methods. For instance, a combination of limited data availability and a restricted computational budget may exacerbate the issue of overfitting. By understanding these complex interactions, we can gain insights into designing more resource-efficient and robust FCL techniques. A detailed analysis will be provided in the next section.
III-B1 Ablation Study on Three Training Resources
Figure 1: Ablation study on three training resources.
1. Does the Memory Buffer matter? As shown in Fig. 1(a), we select {FedAvg, FedCIL, Re-Fed, GLF, CFeD, and AF-FCL} that need extra memory buffer to cache previous samples/synthetic samples/auxiliary datasets. Specifically, when allocated a modest amount of memory buffer, these methods experience a marked improvement in their performance. This enhancement becomes particularly pronounced for methods that leverage generative models and distilled data, particularly when dealing with relatively straightforward datasets characterized by limited variation across data classes and domains. However, as the complexity of the datasets increases, posing challenges such as a wider diversity of classes and domains, the performance of these methods, particularly those reliant on generative modeling, undergoes a noticeable decline. This underscores the limitations of solely relying on synthetic data generation or distillation in complex learning environments. In contrast, methods that employ straightforward sample caching exhibit a more resilient performance profile. They maintain a relatively stable level of performance, demonstrating their robustness against dataset complexity to a certain extent. Nevertheless, as the number of cached samples surpasses a certain threshold, the model’s performance gradually approaches an upper limit, indicating diminishing returns from further sample accumulation. This plateauing effect underscores the need for a delicate balance in managing the memory buffer, aiming to maximize performance gains without incurring unnecessary computational overhead.
2. Does the Computational Budget matter? As depicted in Fig. 1(b), we carefully allocated varying computational budgets for each dataset, taking into account the maximum allowable limit specified in Table III. It is evident from the results that when the computational budgets assigned to the various methods are restricted, the overall performance of the models inevitably suffers. This underlines the critical role that computational resources play in ensuring optimal model performance. However, despite these constraints, certain methods, notably FedAvg and FOT, demonstrate remarkable resilience. They maintain satisfactory performance levels even when operating under limited computational budgets, highlighting their efficiency and effectiveness in resource-constrained environments. On the other hand, when the computational budget is abundant, allowing for unrestricted resource allocation, these other methods display the potential for higher performance ceilings. Given sufficient resources, they may achieve performance levels that rival or surpass those of FedAvg and FOT. Nevertheless, the practical implications of such high resource requirements must be carefully considered, as they may not always be feasible or cost-effective in real-world applications.
3. Does the Label Rate matter? In Fig. 1(c), we delve into the intricate relationship between the label rate and the performance of various model methods. Our analysis reveals a compelling trend: when the label rate falls below 70%, a strategic increase in the proportion of labeled data within the dataset is a potent catalyst for enhancing the performance of most model approaches. However, the Digit-10 dataset stands out as a unique case. Despite boasting a vast quantity of data, the simplicity of its feature information, characterized by single-channel data, renders it less reliant on an extensive labeled dataset for optimal performance. Consequently, even with a relatively modest amount of labeled data, the model can achieve promising results, underscoring the importance of considering the dataset’s inherent characteristics when evaluating the label rate’s impact. This observation underscores the nuanced interplay between the label rate, dataset complexity, and model performance. While a higher label rate generally leads to improved performance, the extent of this improvement can vary significantly depending on the dataset’s characteristics and the model being employed.
Conclusion. Existing methods have encountered limitations in achieving optimal results within resource-constrained environments. Notably, deploying a modestly sized memory buffer has emerged as a crucial factor in significantly augmenting model performance. This enhancement underscores the importance of an adequately provisioned yet economically efficient memory allocation strategy, as an overly generous buffer does not necessarily translate into commensurate gains in performance. A similar trend is observed with the label rate, emphasizing the need for a balanced approach in labeling data to maximize model efficacy. Moreover, the influence of the computational budget on FCL performance is intimately tied to the algorithmic design itself. Simple yet effective methods such as FedAvg demonstrate remarkable resilience, maintaining robust performance even with stringent computational constraints. This finding underscores the potential of streamlined FCL algorithms in addressing real-world challenges posed by limited resources. Experimental evidence highlights the imperative of selecting FCL methods tailored to the specific training resources. Researchers can achieve more effective learning outcomes by aligning algorithmic complexity with the available computational and memory resources.
Figure 2: Analysis of three typical FCL techniques.
III-B2 Analysis of Typical FCL Techniques
1. Does the Sample Caching matter? To gain a deeper understanding of the efficacy of sample caching technology within the constraints of limited training resources, we impose restrictions on the memory buffer size and the label rate illustrated in Fig. 2(a). Regarding the sample caching technique, the label rate of the data plays a pivotal role in determining the quantity and quality of samples available for local training. Specifically, a higher label rate translates into more high-quality, labeled samples within the dataset, significantly enhancing the learning process. The Re-Fed algorithm, in particular, introduces an innovative approach that leverages collaborative storage to intelligently select and retain the most critical samples when both the label rate and memory buffer are limited. This strategy is particularly effective in boosting model performance by focusing on the most informative data points, even under stringent resource constraints. Furthermore, as the label rate and memory buffer size vary, they jointly influence the performance of the sample caching technique in complex ways. The label rate not only dictates the overall quality of the data but also serves as an upper bound on the number of samples that can be feasibly stored within the given memory buffer. Consequently, the effectiveness of the sample caching methodology is intimately tied to the interplay between these two training resources. Optimizing the allocation and utilization of these resources is thus crucial for maximizing the benefits of sample caching under resource-constrained settings.
3. Does the Knowledge Distillation matter? Knowledge distillation is a prevalent technique in mitigating catastrophic forgetting within existing FCL methods. While this approach effectively preserves key information across learning tasks, it necessitates the utilization of additional distillation datasets, whether publicly accessible or synthetically generated. Unfortunately, this requirement introduces additional layers of complexity, including the management of both computational overhead and memory buffer capacity. We primarily focus on limiting the computational budget, as it is often the primary factor constraining the scalability and speed of the distillation process. This emphasis is evident in Fig. 2(c), where the graph clearly illustrates that the computational budget is the primary limiting factor affecting the performance of the corresponding knowledge distillation methods. On the other hand, the memory buffer, though still a consideration, has a relatively minor impact on the overall performance of the knowledge distillation technique. This observation highlights the efficiency of the distillation process, where even a modest quantity of distilled data samples can achieve satisfactory performance gains. By efficiently managing the memory buffer to accommodate these necessary samples, we can maximize the benefits of knowledge distillation without incurring excessive memory overhead. In summary, our approach to knowledge distillation in FCL involves a careful balancing act between computational budget and memory buffer, with a greater emphasis on optimizing the former to ensure the technique’s practicality and performance within the confines of federated and continual learning environments.
Discussion about the Network Extension. In this section, we do not perform additional experimental investigations specifically tailored for the Network Extension method. The main reason is that prevalent network extension methodologies typically presuppose knowledge of task boundaries, implying that the task ID for incoming data is accessible during inference. When this impractical assumption is relaxed, and we analogize to other methodologies that infer the category of the sample without prior task information, these extensions fail to uphold their erstwhile promising performance, as evidenced in Tables II(a) and II(b).
Conclusion. The three prominent FCL strategies can exhibit unique strengths and limitations, particularly concerning the availability of training resources. When resources, including computational power and memory buffer capacity, are abundant, both sample caching and data synthesis methods demonstrate the potential to achieve superior performance. Sample caching, by preserving a subset of representative samples from previous tasks, provides a direct means of mitigating forgetting. Meanwhile, data synthesis, by generating pseudo-data that encapsulates characteristics of previous tasks, offers the advantage of not requiring storage of actual data samples, albeit at the cost of a higher computational burden. This approach relies heavily on a substantial computational budget, as well as a memory buffer, to support data rehearsal, ensuring that the synthesized samples are effectively utilized in refining the model’s knowledge. In contrast, knowledge distillation techniques showcase a remarkable resilience to resource constraints by minimizing the need for extensive memory buffers, as even small amounts of distilled data samples are sufficient to maintain good performance. This characteristic makes knowledge distillation an attractive option for FCL scenarios where resources are limited or constrained. Therefore, selecting the appropriate FCL technique based on training resources is paramount. Sample caching and data synthesis may offer the best path to optimal performance in resource-rich environments. However, in more constrained settings, knowledge distillation emerges as a viable and efficient alternative. Furthermore, further research in this direction is imperative, given the ever-increasing demand for FCL methods that can operate under strict resource limitations. This includes exploring innovative ways to reduce the computational overhead of data synthesis, optimizing memory usage in sample caching strategies, and refining knowledge distillation techniques to make them even more robust and efficient.
IV Quantitative Analysis about two scenarios
TABLE IV: Performance comparison of various methods with a different order of domain tasks. We report the reverse order compared to the Table II(a).
Dataset Metric FedAvg FedCIL Re-Fed FOT
Digit-10 $A(f)$ 75.84±0.36 79.96±1.12 79.63±0.53 76.85±0.74
$\bar{A}$ 83.62±0.41 87.98±0.88 88.23±0.49 85.44±0.83
Office-31 $A(f)$ 36.65±0.52 41.90±1.36 42.71±0.44 42.25±0.79
$\bar{A}$ 41.47±0.49 44.18±0.95 45.05±0.39 45.11±0.62
In this section, we delve deeper into the analysis of existing Federated Continual Learning (FCL) methodologies by conducting extensive quantitative experiments that focus on two crucial aspects: the number of tasks within Class-IL datasets and the order of tasks across various Domain-IL datasets. By systematically varying these parameters, we aim to gain a nuanced understanding of how these factors impact the performance of FCL methods. Firstly, we observe that as the number of tasks in the Class-IL datasets escalates, the performance of all the evaluated methods noticeably declines. Nevertheless, despite this performance decrement, the general trends exhibited by the different methods remain comparable, suggesting that the fundamental challenges they face are similar. Furthermore, our experiments reveal that the order in which domain tasks are presented to the model can significantly influence its ultimate performance within Domain-IL scenarios. This finding emphasizes the importance of task scheduling and curriculum design in FCL systems, as certain sequences may be more conducive to learning than others. However, even with this variability in performance due to task order, each method maintains its distinctive strengths and tendencies, demonstrating its resilience and adaptability to different learning contexts.
V Future Research in Resource-Constrained Federated Continual Learning
@@
  • Sampling Strategy: Sample caching is a mainstream method to prevent catastrophic forgetting in FCL, and it often has good results when memory buffer resources are not tight. In our experiments, we used simple strategies such as Random, FIFO, and LIFO. However, apart from the collaborative replay strategy proposed in [15], there is currently no in-depth exploration of sample caching strategies in FCL. In addition, future research can also focus on directly selecting exemplar samples during local training to train and cache locally.
@@
VI Conclusion
This paper has presented a tutorial on FCL under resource constraints. Firstly, we begin with an introduction to the existing FCL methods and three training resources on distributed devices that influence the performance of methods. Then, we conduct experiments using ablation studies on each training resource and analyze the effect of four typical FCL techniques on alleviating catastrophic forgetting in FCL. Finally, we discuss future research directions in resource-constrained FCL.

### v2

@@ B Problem Formulation
Table 5: Experimental Details. Analysis of various considered settings of different datasets in the experiments section.
@@
[…]
@@
B.1 Experiment Setup
B.1.1 Datasets.
Our experiments utilize diverse datasets distributed across two federated incremental scenarios, encompassing six datasets in total.
Class-Incremental Task Dataset: This dataset gradually introduces new classes. It begins with a subset of classes and expands by adding more classes in subsequent stages, facilitating models to learn and accommodate an increasing range of classes.
(1) CIFAR-10 krizhevsky2009learning : Comprises 10 object classes, encompassing everyday items, animals, and vehicles. It includes 50,000 training and 10,000 test images.
(2) CIFAR-100 krizhevsky2009learning : Similar to CIFAR-10 but with 100 detailed object classes. It contains 50,000 training and 10,000 test images.
(3) Tiny-ImageNet le2015tiny : A selection from the ImageNet dataset, featuring 200 object classes. It includes 100,000 training images, 10,000 validation images, and 10,000 test images.
Domain-Incremental Task Dataset: This dataset progressively introduces new domains. It starts with samples from a specific domain and incorporates additional domains in later stages, allowing models to generalize to new contexts.
(1) Digit-10: The Digit-10 dataset includes 10 digit categories across four domains: MNIST lecun2010mnist , EMNIST cohen2017emnist , USPS hull1994database , and SVHN netzer2011reading . Each dataset represents a specific domain, such as handwriting style, and contains 10 digit classes.
  • MNIST: Contains 60,000 training and 10,000 test handwritten digit images.
  • EMNIST: An extension of MNIST with 240,000 training and 40,000 test handwritten character (letter and digit) images.
  • USPS: Includes 7,291 training and 2,007 test handwritten digit images from the US Postal Service.
  • SVHN: Features 73,257 training and 26,032 test images of house numbers captured from Google Street View.
(2) Office-31 saenko2010adapting : Contains images from three domains: Amazon, Webcam, and DSLR. It covers 31 categories, with approximately 4,100 images per domain.
(3) Office-Caltech-10 zhang2020impact : Includes images from four domains: Amazon, Caltech, Webcam, and DSLR. It comprises 10 object categories, with around 2,500 images per domain.
B.1.2 Baselines.
In this paper, we follow the same protocols proposed by mcmahan2017communication ; rebuffi2017icarl to set up tasks. We evaluate our resource-constrained settings with the following baselines.
FedAvg mcmahan2017communication : It is a representative FL model that aggregates client parameters in each communication round. It is a simple but effective model for FL.
FedProx li2020federated : It is also a representative FL model, which is better at handling heterogeneity in federated networks than FedAvg.
FedAvg/FedProx+ER rolnick2019experience : Experience Replay is a technique primarily utilized in reinforcement learning, especially in deep reinforcement learning algorithms. Here, we combine the ER algorithm with the FL algorithms FedAvg and FedProx.
GLFC dong2022federated : This method addresses the federated class-incremental learning issue and trains a global model with additional class-imbalance losses. A proxy server is used to reconstruct samples to help select the best old models for model updates.
FedCIL qi2023better : This method employs the ACGAN network to generate synthetic samples to consolidate the global model and align sample features in the output layer. The authors conducted experiments in the FCIL scenario, and we adopted it in our FDIL setting.
TARGET zhang2023target : This method uses a generator to synthesize data to simulate the global data distribution on each client without additional data from previous tasks. The data heterogeneity is considered with catastrophic forgetting, and the authors focus on the FCIL scenario.
AF-FCL wuerkaixi2023accurate : This method focuses on the data noise among previous tasks, which degrades the model performance. The NF model is used to quantify the credibility of previous knowledge and select the transfer knowledge.
Re-Fed li2024efficient : This approach proposes synergistic replay for each client to selectively cache samples based on the understanding of both local and global data distributions. Each client trains an additional personalized model to discern the importance of the score to each client.
FOT bakman2023federated : This method uses orthogonal projection technology to project different parameters into different spaces to isolate task parameters. In addition, this method proposes a secure parameter aggregation method based on projection. In the model inference stage, the method requires assuming that the boundaries of the task are known. We modified it to an automated inference of the model.
CFeD ijcai2022p0303 : It employs public datasets as proxy data for knowledge distillation at both client and server sides. Upon introducing new tasks, clients use proxy data to transfer knowledge from previous to current models thus mitigating inter-task forgetting.
FedWeIT yoon2021federated : This method divides parameters into task-specific and shared parameters. A multi-head model based on known task boundaries is used in the original method. To ensure a fair comparison, we modify it to automate the inference of the model.
C Future Research in Resource-Constrained Federated Continual Learning
@@
  • Sampling Strategy: Sample caching is a mainstream method to prevent catastrophic forgetting in FCL, and it often has good results when memory buffer resources are not tight. In our experiments, we used simple strategies such as Random, FIFO, and LIFO. However, apart from the collaborative replay strategy proposed in li2024efficient , there is currently no in-depth exploration of sample caching strategies in FCL. In addition, future research can also focus on directly selecting exemplar samples during local training to train and cache locally.